OpenStudy (anonymous):

what is the solution of l2x-1 -8 =18 ?

OpenStudy (anonymous):

what is that in the left side of 2?

OpenStudy (jhannybean):

First of all,you want to isolate the x on one side of the equation and put all the other numbers pn the other side, so we start by getting rid of the numbers can you tell me what -1 - 8 =?

OpenStudy (anonymous):

nevamind wrong equation 2l 2x-1 l -8=18?

OpenStudy (anonymous):

^ that's the one im confused on

OpenStudy (jhannybean):

ok what is -11 -8?

OpenStudy (anonymous):

theres no negative 11 there those line tingys

OpenStudy (anonymous):

2I 2x-2I=26 this is a mod function. the speciality in it is that when you take off mod, you have to consider both positive and negative values. =>I 2x-2I=13 =>2x-2=(+-)13 you have to do it separately. 2x-2=13 =>x=15/2 and when you consider negative sign, 2x-2=-13 x=-(11/2)